OpenStudy (anonymous):

The giant monster NADIAZILLA is dangling Doug by his belt. Doug weighs 650 N. If NADIAZILLA accelerates Doug upwards at 3.5 m/s^2 before flinging him into the Ottawa River, what is the tension in Doug's belt?

OpenStudy (anonymous):

The tension in Doug's belt will be equal to the applied force (the force that the moster exerts on Doug). F(weight) = 650 N a = 3.5 m/s^2 g = 9.81 m/s^2 F(weight) = mg m = 650/9.81 = 66.26 kg F(net) = ma F(net) = (66.26)(3.5) = 232 N F(net) = F(applied) - F(weight) F(applied) = F(net) + F(weight) F(applied) = 232 + 650 = 882 N Asnwer: 882 N